OpenStudy (anonymous):

A conveyor belt carries supplies from the first floor to the second floor, which is 26 feet higher. The belt makes a 60° angle with the ground. How far do the supplies travel from one end of the conveyor belt to the other? Round your answer to the nearest foot. If the belt moves at 75 ft/min, how long, to the nearest tenth of a minute, does it take the supplies to move to the second floor? A. 45 ft; 37 min B. 37 ft; 22.5 min C. 30 ft; 0.4 min D. 15 ft; 1.1 min Reset Selection

OpenStudy (anonymous):

HELP PLEASE IM ON A TIMED EXAM

OpenStudy (radar):

 sin 60 = 26/x x=26/sin60 x=30 ft

OpenStudy (radar):

Looks like Option C is the best choice.